OpenStudy (anonymous):

The sum of two consecutive integers is at least 46. What is the least possible pair of integers? (Points : 1) 21 and 23 23 and 24 22 and 23 24 and 25

OpenStudy (anonymous):

@Euler271 do u know how to do this problem or r u just viewing

OpenStudy (anonymous):

lol i opened many tabs. i know how to do it. havent gotten to it yet lol

OpenStudy (anonymous):

oh just asking lol @Euler271

OpenStudy (anonymous):

2 consecutive integers at least 46: \[x + (x+1) \ge 46\] solve for x

OpenStudy (anonymous):

\[2x \ge 45\]\[x \ge 22.5\] so the smaller integer is greater than 22.5 and the second one is greater than 23.5. only one of the four options works

OpenStudy (anonymous):

actually 2 options work. you must pick the one with the smallest integers

OpenStudy (anonymous):

23 and 24